Hello,
I need help with the following math problem:
Exits on a highway are numbered consecutively from 1 to 50. The distance from exit 41 to exit 50 is 100 km. If each exit is at least 6 km from the next exit, what is the longest possible distance, in kilometers, between exit 47 and exit 48?
Thanks,
Nicole

Answer by ikleyn(52778)

.
Hello,
I need help with the following math problem:
Exits on a highway are numbered consecutively from 1 to 50. The distance from exit 41 to exit 50 is 100 km.
If each exit is at least 6 km from the next exit, what is the longest possible distance, in kilometers, between exit 47 and exit 48?
Thanks,
Nicole
~~~~~~~~~~~~~~~~~~~~~~~

Assume that the distance between each other pair of exits is exactly 6 km. It doesn't contradict to the condition, hence, it is allowed.

The total number of intervals in the highway between exits 41 and 50 is 9 (nine).

The number of intervals others than (47-48) is 9-1 = 8.

8*6 = 48.

100 - 48 = 52.

Answer.  the longest possible distance, in kilometers, between exit 47 and exit 48  is 52 kilometers.